The fast server in women's tennis is Venus Williams, who recorded a serve of 130 mi/h (209 km/h) in 2007.

If her racket pushed on the ball for a distance of 0.11 m, what was the average acceleration of the ball during her serve?
Answer #1

Here ,

speed of shot , u = 130 mi/hr

u = 209 km/hr = 209 * 1000/3600 m/s

u = 58.1 m/s

distance , d = 0.11 m

let the average acceleration is a

Using third equation of motion

v^2 - u^2 = 2 * a * d

58.1^2 = 2 * a * 0.11

a = 15343.7 m/s^2

the acceleration of the ball is 15343.7 m/s^2

b)

let the time interval is t

Using second equation of motion

s = u * t + 0.5 * a * t^2

0.11 = 0.5 * 15343.7 * t^2

t = 3.78 *10^-3 s

the time interval of the contact is 3.78 *10^-3 s